You form a portfolio of stocks. Stock A has a beta of 0.9, Stock B has a beta of 1.8, and Stock C has a beta of 1.3. If 13% of your portfolio is invested in stock A, and 45% of your stock is invested in Stock B, what is your portfolio beta?

(Enter your response to two decimal places ex: 1.23)
Answer #1

Investment in C=100-(13+45)=42%

Portfolio beta=Respective beta*Respective weight

=(0.13*0.9)+(0.45*1.8)+(0.42*1.3)

which is equal to

=1.47(Approx).